A typist makes an average of 5 typo errors per page that she types. If the number of errors per each typed page can be modeled according to the Poisson distribution, what is the probability that she will make more than 3 typo errors in the next page?
Answer #1

Average = 5 type errors per page

Hence,

P(More than 3 typo)

= 1 - P(Less than or equal to 3 typo)

= 1 - [P(0) + P(1) + P(2) + P(3)]

= 1 - [\frac{e^{-5}5^0}{0!} + \frac{e^{-5}5^1}{1!} + \frac{e^{-5}5^2}{2!} + \frac{e^{-5}5^3}{3!}]

= 0.7350